Solving by Elimination.
{x-2y=14
3x-y=2

Answer by mananth(16949)

x-2y=14 .............1
3x-y=2 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -2
x-2y=14
-6x+2y=-4
Add the two equations
-5x=10
/-5
x=-2
plug value of x in (1)
x-2y=14
-2-2y=14
-2y=14+2
-2y=16
y=-8
m.ananth@hotmail.ca
CHECK

1 x + -2 y = 14
-2 + 16 = 14